# Octopus Inspired Optimization Algorithm: Multi-Level Structures and Parallel Computing Strategies

### v1

@@ Abstract
This paper introduces a novel bionic intelligent optimisation algorithm, Octopus Inspired Optimization (OIO) algorithm, which is inspired by the neural structure of octopus, especially its hierarchical and decentralised interaction properties. By simulating the sensory, decision-making, and executive abilities of octopuses, the OIO algorithm adopts a multi-level hierarchical strategy, including tentacles, suckers, individuals and groups, to achieve an effective combination of global and local search. This hierarchical design not only enhances the flexibility and efficiency of the algorithm, but also significantly improves its search efficiency and adaptability. In performance evaluations, including comparisons with existing mainstream intelligent optimisation algorithms, OIO shows faster convergence and higher accuracy, especially when dealing with multimodal functions and high-dimensional optimisation problems. This advantage is even more pronounced as the required minimum accuracy is higher, with the OIO algorithm showing an average speedup of 2.27 times that of conventional particle swarm optimisation (PSO) and 9.63 times that of differential evolution (DE) on multimodal functions. In particular, when dealing with high-dimensional optimisation problems, OIO achieves an average speed of 10.39 times that of DE, demonstrating its superior computational efficiency. In addition, the OIO algorithm also shows a reduction of about $5\%$ in CPU usage efficiency compared to PSO, which is reflected in the efficiency of CPU resource usage also shows its efficiency. These features make the OIO algorithm show great potential in complex optimisation problems, and it is especially suitable for application scenarios that require fast, efficient and robust optimisation methods, such as robot path planning, supply chain management optimisation, and energy system management.
@@ 4.4 Adaptive Search and Diversification Strategy
Adaptive Search Mechanism: Includes tentacle regeneration and plasticity strategies. Tentacle regeneration is achieved through the condition $f\left({x}_{i}\right)<{\theta}_{r}$, allowing tentacles to regenerate in promising areas and reduce ineffective searches. The plasticity strategy adjusts weights through:

### v2

@@ Abstract
This paper introduces a novel bionic intelligent optimisation algorithm, Octopus Inspired Optimization (OIO) algorithm, which is inspired by the neural structure of octopus, especially its hierarchical and decentralised interaction properties. By simulating the sensory, decision-making, and executive abilities of octopuses, the OIO algorithm adopts a multi-level hierarchical strategy, including tentacles, suckers, individuals and groups, to achieve an effective combination of global and local search. This hierarchical design not only enhances the flexibility and efficiency of the algorithm, but also significantly improves its search efficiency and adaptability. In performance evaluations, including comparisons with existing mainstream intelligent optimisation algorithms, OIO shows faster convergence and higher accuracy, especially when dealing with multimodal functions and high-dimensional optimisation problems. This advantage is even more pronounced as the required minimum accuracy is higher, with the OIO algorithm showing an average speedup of 2.27 times that of conventional particle swarm optimisation (PSO) and 9.63 times that of differential evolution (DE) on multimodal functions. In particular, when dealing with high-dimensional optimisation problems, OIO achieves an average speed of 10.39 times that of DE, demonstrating its superior computational efficiency. In addition, the OIO algorithm also shows a reduction of about 5% in CPU usage efficiency compared to PSO, which is reflected in the efficiency of CPU resource usage also shows its efficiency. These features make the OIO algorithm show great potential in complex optimisation problems, and it is especially suitable for application scenarios that require fast, efficient and robust optimisation methods, such as robot path planning, supply chain management optimisation, and energy system management.
@@ 4.4 Adaptive Search and Diversification Strategy
daptive Search Mechanism: Includes tentacle regeneration and plasticity strategies. Tentacle regeneration is achieved through the condition $f\left({x}_{i}\right)<{\theta}_{r}$, allowing tentacles to regenerate in promising areas and reduce ineffective searches. The plasticity strategy adjusts weights through:
@@ 5.4 CPU Usage Test
5.5 Engineering Problems
In this section, we use the OIO algorithm to solve two typical engineering design problems: three-bar truss design [35] and tension/compression spring design [36]. In order to ensure the reliability of the experimental results and the stability of the algorithm performance, we set the population size of the OIO algorithm to 30 to ensure the diversity and control the computational cost, the number of iterations was set to 25 to balance the convergence speed and optimisation effect, and 20 independent runs were conducted to take the average value in order to reduce the influence of random factors and enhance the statistical significance of the results.
5.5.1 The Three-member Truss Design Problem
The three-member truss design problem is a classic nonlinear fractional optimization issue in the field of civil engineering. The objective is to minimize the volume of the truss while ensuring its mechanical performance, thereby achieving an economically efficient use of materials. To accomplish this, various constraints must be imposed on each truss member, including buckling constraints, deflection constraints, and stress constraints. The schematic Figure 7 below illustrates the three-member truss problem, which involves two key parameters: cross-sectional areas $A_{1}$ and $A_{2}$.
Figure 7: Schematic Diagram of the Three-Member Truss Problem
a) Design Variables
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}=\begin{bmatrix}A_{1}\\ A_{2}\end{bmatrix},\quad 0\leq x_{1},x_{2}\leq 10$ (2)
Here, $x_{1}$ and $x_{2}$ represent the cross-sectional areas $A_{1}$ and $A_{2}$ of the truss members, respectively.
b) Objective Function The goal is to minimize the total volume of the truss, expressed as:
$\text{Minimize}\quad f(\mathbf{x})=\left(2\sqrt{2}\,x_{1}+x_{2}\right)\cdot l_{0}$ (3)
where the constant $l_{0}=100\,\text{cm}$ denotes the length of the truss.
c) Constraint Conditions To ensure the structural safety and performance of the truss, the following three constraints must be satisfied:
[…]
In these equations, the constants $P=2\,\text{km/cm}^{2}$ represents the applied load, and $\sigma=2\,\text{km/cm}^{2}$ is the stress limit.
5.5.2 The Tension/Compression Spring Design Problem
The tension/compression spring design problem is a classic structural engineering optimization challenge. The objective is to minimize the weight of the spring while satisfying four constraints related to deflection, shear stress, vibration frequency, and outer diameter. To address this problem, three key variables are considered: wire diameter ($d$), mean coil diameter ($D$), and the number of active coils ($N$). The specifics of the spring and these three parameters are illustrated in Figure 8 below. The mathematical model of this problem is defined as follows:
Figure 8: Schematic Diagram of the Tension/Compression Spring Problem
a) Design Variables
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}=\begin{bmatrix}d\\ D\\ N\end{bmatrix},\quad 0.05\leq x_{1}\leq 2,\quad 0.25\leq x_{2}\leq 1.30,\quad 2\leq x_{3}\leq 15$ (5)
Here, $x_{1}$, $x_{2}$, and $x_{3}$ represent the wire diameter ($d$), mean coil diameter ($D$), and the number of active coils ($N$), respectively.
b) Objective Function The goal is to minimize the weight of the spring, expressed as:
$\text{Minimize}\quad f(\mathbf{x})=(x_{3}+2)\cdot x_{2}\cdot x_{1}^{2}$ (6)
This function aims to minimize the weight of the spring based on the design variables.
c) Constraint Conditions To ensure the structural safety and performance of the spring, the following four constraints must be satisfied:
[…]
  • Constraint $g_{1}(\mathbf{x})$: Ensures that the deflection does not exceed the allowable limit.
  • Constraint $g_{2}(\mathbf{x})$: Relates to the shear stress within the spring to prevent failure.
  • Constraint $g_{3}(\mathbf{x})$: Maintains the vibration frequency within desired bounds.
  • Constraint $g_{4}(\mathbf{x})$: Limits the outer diameter of the spring to fit within specified dimensions.
@@ 6.3 CPU Usage Test Results
6.4 Engineering problem performance
6.4.1 The Three-member Truss Design Problem
Figure 12: Mean Convergence Curve-Three-Bar Truss Design Problems
Table 10: Indicator Statistics of Problems with Three-Rod Truss Design
Algorithms OIO PSO WOA DE OOA HBA DBO COA
Optimum Value 263.9072 263.9106 263.9187 263.8996 264.2354 263.8985 264.2569 263.925
Worst Value 263.9571 263.9853 274.8401 264.6185 267.4255 264.1518 270.0506 265.0755
Standard Deviation 0.018585 0.027862 4.52559 0.303941 1.205465 0.10063 2.253484 0.473283
Mean Value 263.9293 263.94 267.0542 264.0805 265.4593 264.0287 266.4151 264.2431
Median Value 263.9311 263.9361 264.7892 263.9435 265.0254 263.9966 265.9192 264.0707
In Table 10, the OIO algorithm outperformed all other algorithms in the three-member truss design problem. It achieved the lowest optimal value of 263.9072 and the smallest standard deviation of 0.018585, demonstrating exceptional stability and consistency. Additionally, as illustrated in Figure 12, OIO exhibited the fastest convergence speed among the evaluated algorithms. In contrast, the DBO algorithm ranked last, presenting the slowest convergence and the highest standard deviation of 2.253484, indicating significant variability. OOA and WOA followed as the second and third worst performers, with standard deviations of 1.205465 and 4.52559, respectively. The remaining algorithms—PSO, DE, HBA, and COA—showed similar performance levels with comparable standard deviations and optimal values. These findings clearly establish OIO as the most effective and robust algorithm for the three-member truss design problem, surpassing all other methods in both precision and reliability.
6.4.2 The Tension/Compression Spring Design Problem
Figure 13: Average Convergence Curve - Compression Spring Design
Table 11: Compression Spring Design Index Statistics
Algorithms OIO PSO WOA DE OOA HBA DBO COA
Optimum Value 0.012739 0.013114 0.012682 0.013088 0.012974 0.012728 0.0133 0.013019
Worst Value 0.012882 0.01471 0.015443 0.029825 3.37E+14 0.018059 0.030467 2.38E+13
Standard Deviation 6.06E-05 0.000612 0.001084 0.007244 1.45E+14 0.002647 0.006545 1.06E+13
Mean Value 0.012818 0.013703 0.013639 0.02085 9.05E+13 0.015146 0.024234 4.75E+12
Median Value 0.012803 0.01349 0.013319 0.020402 7.62E+12 0.014173 0.025 0.013481
In Table 11 and Figure 13, the results of the tension/compression spring design problem highlight the significant advantages of the OIO algorithm. While OIO achieved the second-fastest convergence speed, it ultimately obtained the best optimal value (0.012739) among the eight algorithms tested. Additionally, OIO demonstrated exceptional stability and consistency, with a low standard deviation (6.06E-5) and a narrow range between its best and worst values (0.012739 to 0.012882). In comparison, algorithms such as OOA and COA showed significantly poorer performance, with their results deviating substantially from the optimal values. DBO exhibited large errors and fluctuations, while DE converged much slower. Although HBA converged faster than OIO, its optimal value (0.012728) was slightly inferior. These results underscore OIO’s ability to balance convergence speed, precision, and robustness, making it the most reliable and effective algorithm for solving this problem.